OpenStudy (anonymous):

solve the nonlinear inequality: 3\(x-1) - 4\x > or = 1

OpenStudy (anonymous):

'\' mean division,right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay first take L.C.M it will give you [3x-4(x-1)]/x(x-1)>=1 then it will give you by simplification : (1-x)>=x^2-x this implies 1>=x^2 take under rrot on both sides you will get x<=1

OpenStudy (anonymous):

can you show the steps of simplification please ?

OpenStudy (anonymous):

(3/x-1)-(4/x)>=1 [3(x)-4(x-1)/x(x-)]>=1 (L.C.M) [1-x/x(x-1)]>=1 1-x>=x(x-1) (Multiply both sides by x(x-) ) 1-x>=x^2-x 1-x+(x)>=x^2-x+(x) (Adding x on both sides) 1>=x^2 taking sqrt on both sides 1>=x this implies x<=1 OKAY NOW?

OpenStudy (anonymous):

this answer does not appear to work when inserted in the equation

OpenStudy (anonymous):

x must not be equal to 0 and 1 because they make this undefined.Everthing else works